# An alternative theory of magnetic flux tubes in strong fields via axion origin photons

### v2

@@ Abstract
In our alternative theory, built around the coincidence of experimental and theoretical data, three ”free” parameters – the magnetic field in the tachocline of the order of $\sim 10^{7}~G$ (see Fig. (A.1) and Eq. (A17) in V. D. Rusov et al. (2021)), the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Eq. (11) in V. D. Rusov et al. (2021)), and the asymmetric dark matter (ADM) in the Universe with $m_{ADM}\sim 5~GeV$ ((see V. D. Rusov et al. (2021); A. C. Vincent et al. (2016)) – give a complete solution to the problem of the theory of magnetic flux tubes in strong fields with 11-year variations of axion-origin photons, which are caused by and anticorrelated to the 11-year variations in density of ADM, gravitationally captured on the Sun.
[…]
show]siiis@te.net.ua
@@ I Introduction
It is known that the unsolved problem of energy transport by magnetic flux tubes at the same time represents another unsolved problem related to the sunspot darkness (see 2.2 in M. Rempel & R. Schlichenmaier (2011)). Of all the known concepts playing a noticeable role in understanding the connection between the energy transport and sunspot darkness, let us consider the most significant theory, in our view. It is based on the Parker-Biermann cooling effect (E. N. Parker, 1955a; L. Biermann, 1941; E. N. Parker, 1979a) and originates from the early works of L. Biermann (1941) and H. Alfvén (1942).
@@
As you know, the Parker-Biermann cooling effect (E. N. Parker, 1955a; L. Biermann, 1941; E. N. Parker, 1979a), which plays a role in our current understanding, originates from L. Biermann (1941) and H. Alfvén (1942): in a highly ionized plasma, the electrical conductivity can be so large that the magnetic fields are frozen into the plasma. Biermann realized that the magnetic field in the spots themselves can be the cause of their coolness – it is colder because the magnetic field suppresses the convective heat transport. Hence, the darkness of the spot is due to a decrease in surface brightness.
@@
Parker (E. N. Parker, 1955a, 1974a, 1974b, 1974, 1979a) has pointed out that the magnetic field can be compressed to the enormous intensity only by reducing the gas pressure within the flux tube relative to the pressure outside, so that the external pressure compresses the field. The only known mechanism for reducing the internal pressure sufficiently is a reduction of the internal temperature over several scale heights so that the gravitational field of the Sun pulls the gas down out of the tube (as described by the known barometric law $dp/dz=-\rho g$). Hence it appears that the intense magnetic field of the sunspot is a direct consequence of the observed reduced temperature (E. N. Parker, 1955a).
@@
On the other hand, Parker (E. N. Parker, 1974, 1977) has also pointed out that the magnetic inhibition of convective heat transport beneath the sunspot, with the associated heat accumulation below, raises the temperature in the lower part of the field. The barometric equilibrium leads to enhanced gas pressure upward along the magnetic field, causing the field to disperse rather than intensify. Consequently, E. N. Parker (1974) argued that the temperature of the gas must be influenced by something more than the inhibition of heat transport!
@@
Our unique alternative idea is that the explanation of sunspots is based not only on the suppression of convective heat transfer by a strong magnetic field (of the order of $\sim 10^{7}~G$ (V. D. Rusov et al., 2021)) through the enhanced cooling of the Parker-Biermann effect (E. N. Parker, 1974a), but also on the appearance of the axions of photonic origin (Fig 3, Fig. B.1 in V. D. Rusov et al. (2021)) from the tachocline to the photosphere, which is confirmed by the “disappearance” of the heat, and consequently, the temperature in the lower part of the magnetic tube (E. N. Parker, 1974, 1977) due to the axions of photonic origin from the photon-axion oscillations in the O-loop near the tachocline (see Fig. 3).
@@
This means that the appearance of axions of photonic origin, which remove the problem of the temperature rise in the lower part of the magnetic tube, and the photons of axionic origin, which have the free path (Rosseland length; see Fig. B.3 in (V. D. Rusov et al., 2021)) from the tachocline to the photosphere, are the explanation of sunspots based not only on the suppression of convective heat transport by a strong magnetic field (V. D. Rusov et al., 2021), but also on the indispensable existence of the Parker-Biermann cooling effect. At the same time, we clearly understand that stronger fields would seriously suppress the action of the dynamo (see (E. E. Deluca & P. A. Gilman, 1986)).
@@
On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in V. D. Rusov et al. (2021)). Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect (A. V. Ettingshausen & W. Nernst, 1886; E. H. Sondheimer, 1948; L. J. Spitzer, 1956; Y. B. Kim & M. J. Stephen, 1969) creates electric currents that are inversely proportional to the strong magnetic field of the tachocline. As we showed earlier (V. D. Rusov et al., 2021), the toroidal magnetic field of tachocline by means of the thermomagnetic Ettingshausen-Nernst effect
@@
Using the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in V. D. Rusov et al. (2021)).
Since the variations of axion origin photons are the result of the changes in Sun luminosity, then unlike a self-excited dynamo, an unexpected yet simple question arises: is there a dark matter chronometer hidden deep in the Sun? And consequently, the next question arises: who controls the Sun or who measures the precise “clock” inside the Sun’s core?
Figure 1: An illustration of the main possible processes of a magnetically active star of the Sun type. (a) $\alpha$-effect, $\Omega$-effect and BL mechanism as components of the solar dynamo model. The $\Omega$-effect (blue) depicts the transformation of the primary poloidal field into a toroidal field by differential rotation. Regeneration of the poloidal field is then performed either by the $\alpha$-effect (top) or by the BL mechanism (yellow in the middle). In case of $\alpha$-effect, the toroidal field at the base of the convection zone is subject to cyclonic turbulence. In the BL mechanism, the main process of regeneration of the poloidal field (based on the $\Omega$-effect (blue)) is the formation of sunspots on the surface of the Sun from the rise of buoyant toroidal flux tubes from the base of the convection zone. The magnetic fields of these sunspots closest to the equator in each hemisphere diffuse and join, and the field due to the spots closer to the poles has a polarity opposite to the current that initiates rotation of the polarity. The newly formed polar magnetic flux is transported by the meridional flow to deeper layers of the convection zone, thereby creating a new large-scale poloidal field. Derived from S. Sanchez et al. (2014). (b) 11-year variations of sunspot magnetic fields as a component of the solar anti-dynamo. The formation of the thermomagnetic EN effect (see L. J. Spitzer (1962, 2006); V. Rusov et al. (2015)), emphasizes that this process is associated with the continuous transformation of toroidal field into poloidal one ($T\rightarrow P$ transformation), but not vice versa ($P\rightarrow T$). For this purpose, we note (seeV. D. Rusov et al. (2021)) that the modulations of the ADM halo in the Galactic Center (GC), which is an indicator of the periods of S-stars, e.g. S-102 with a period of about 11 years, lead (via vertical density waves from the black hole to the solar neighborhood) to modulations of the ADM density, gravitationally trapped in the interior of the Sun. This means that anticorrelation modulations of the ADM density control the solar luminosity or sunspot cycles, in which the sunspot variations are identical to the 11-year variations of the magnetic flux tubes (see Section II): maximum (red lines) and minimum (green-red lines).
In order to answer this question, let us first consider all the unexpected and intriguing implications of the 11-year modulations of the ADM density in the solar interior and around the BH (see Sect. 3 in V. D. Rusov et al. (2021)).
A unique result of our model is the fact (see Sect. 3 in V. D. Rusov et al. (2021)), that the periods, velocities and modulations of the S-stars are the essential indicator of the modulation of the ADM halo density in the fundamental plane of the Galaxy center, which closely correlates with the density modulation of the baryon matter near the supermassive BH. If the modulations of the ADM halo at the GC lead to modulations of the ADM density on the surface of the Sun (through vertical density waves from the disk to the solar neighborhood), then there is an ”experimental” anticorrelation identity between such indicators as the ADM density modulation in the solar interior and the number of sunspots. Or equivalently, between the modulation of solar axions (or photons of axion origin) and the sunspot cycles!
@@
A hypothetical pseudoscalar particle called axion is predicted by the theory related to solving the CP-invariance violation problem in QCD. The most important parameter determining the axion properties is the energy scale $f_{a}$ of the so-called U(1) Peccei-Quinn symmetry violation. It determines both the axion mass and the strength of its coupling to fermions and gauge bosons including photons. However, in spite of the numerous direct experiments, axions have not been discovered so far. Meanwhile, these experiments together with the astrophysical and cosmological limitations leave a rather narrow band for the permissible parameters of invisible axion (e.g. $10^{-6}eV\leqslant m_{a}\leqslant 10^{-2}eV$ (G. G. Raffelt, 2004, 2008a)). The PQ mechanism, solving the strong CP problem in a very elegant way (R. D. Peccei & H. R. Quinn, 1977a, b; F. Wilczek, 1978; S. Weinberg, 1978), is especially attractive here, since the axion is also a candidate for dark matter (J. Preskill et al., 1983; L. Abbott & P. Sikivie, 1983; M. Dine & W. Fischler, 1983; M. Kawasaki & K. Nakayama, 2013; D. J. Marsh, 2016; L. Di Luzio et al., 2020; P. Sikivie, 2021).
@@
At present, we have shown that axions born in the core of the Sun, and photons of axion origin (in the solar corona with a temperature of one million degrees) yield a tight constraint on the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ and mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Fig. 2a; also Fig.1 in V. D. Rusov et al. (2021)). Most importantly, hadronic axions produced in the solar core are controlled by anticorrelated 11-year variations in the density of asymmetric dark matter (ADM) gravitationally trapped in the solar interior (via vertical density waves from the black hole disk to the solar neighborhood), with the hard part of the solar photon spectrum being of axion origin, which is defined as the product of the fraction of axions in the solar core and the fraction of the sunspot area. The latter is the result of solar luminosity variations (see Eqs. (27)-(28) in V. D. Rusov et al. (2021)). This means that solar luminosity variations and other solar cycles are the result of 11-year variations in ADM density, which are anticorrelated with the density of hadronic axions in the solar core.
Figure 2: (a) Summary of astrophysical, cosmological and laboratory constraints on axions. Comprehensive axion parameter space, highlighting two main front lines of direct detection experiments: helioscopes (CAST (CAST Collaboration, 2017)) and haloscopes (ADMX (S. J. Asztalos et al., 2010), and RBF (W. U. Wuensch et al., 1989) and microwave resonators (B. M. Brubaker et al., 2017; L. Zhong et al., 2018; T. Braine et al., 2020; S. Lee et al., 2020; K. M. Backes et al., 2021)). The astrophysical bounds from horizontal branch and massive stars are labeled “HB” (G. G. Raffelt, 2008b) and “Cepheids” (G. Carosi et al., 2013), respectively, and there are also astrophysical hints (WD cooling hints, and HB hint). The QCD motivated models (KSVZ (J. E. Kim, 1979; M. Shifman et al., 1980)) for axions lay in the yellow diagonal band. A plot of $g_{a\gamma}$ versus $m_{a}$ with the most stringent results (solid lines) and sensitivity perspectives (dashed lines) of observations and experiments directly comparable to the different phases of IAXO are shown, BabyIAXO, IAXO, and an upgraded version of IAXO, IAXO+. The yellow band denotes the region of the parameter space favoured by QCD axion models. The red star marks the values of the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ and the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, which were first obtained experimentally (see V. D. Rusov et al. (2021)).
(b) $R$ parameter constraints, which compares the numbers of stars in the the horizontal branch (HB) and in the upper portion of the red giant branch (RGB), to helium mass fraction $Y$ and axion coupling $g_{a\gamma}$ (adopted from (A. Ayala et al., 2014)). The resulting bound on the axion ($g_{10}=g_{a\gamma}/(10^{-10}~GeV^{-1})$ is somewhere between rather conservative $0.5<g_{10}<0.8$ and most aggressive $0.35<g_{10}<0.5$ (A. Friedland et al., 2013). The red line marks the value of the axion–photon coupling constant $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$ adopted from Eq. (2) in (A. Ayala et al., 2014). The blue shaded area represents the bounds from Cepheids observation. The yellow star corresponds to $Y=0.254$ and the bounds from HB lifetime (yellow dashed line).
@@
Interestingly, in contrast to our Fig. 2a, in Fig. 2b, we can see a well-known strong constraint on the axion-photon coupling $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$, but in a wide mass range (see Fig. 2b and Eq. (2) in (A. Ayala et al., 2014)) from an analysis of a sample of 39 Galactic Globular Clusters. However, in our case (Fig. 2a) the wide range of masses in (A. Ayala et al., 2014) iso, ddly enough, replaced by a single mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, which fits not only the axions of galactic globular clusters (A. Ayala et al., 2014), but also all axion stars in our galaxy, including neutron stars (M. Buschmann et al., 2021), supernovae (P. Carenza et al., 2019, 2021), and the Sun (V. D. Rusov et al., 2021).
@@
II Physics of the theory of practically empty magnetic flux tubes at strong fields, $\sim 10^{7}~G$, at which there is a free path for photons of axion origin from the tachocline to the photosphere.
This raises the fundamental question of the self-exciting action of the dynamo, the remarkable physics of which was discovered some 70 years ago (see E. N. Parker (1955b, 1957)), beginning with the advent of Babcock’s solar magnetograph (H. W. Babcock & H. D. Babcock, 1955), and still today, but which, as no surprisingly, has not received a complete and final solution. Avoiding an alternative theory, most physicists believe that it is necessary to study more deeply the known physical problems of various aspects of the theory of solar dynamo (see B. B. Karak et al. (2014)) before passing a verdict on it. This continues to this day (see B. B. Karak (2024)).
In order for our alternative theory (for example, magnetic flux tubes in strong fields using photons of axion origin) to solve all its problems (see Sections II.1-II.2), it is necessary to solve the most important problem: In what way come sunspots, generated by a strong magnetic field in a flux tube from the base of the convection zone to the surface of the Sun, have 11-year variations without any dynamo?
The answer is very simple! Almost all physicists, and in particular dynamo physicists, know that in the Universe, and, of course, in our galaxy, there are ADM, and the highest density of ADM is located exactly at the center of the Galaxy, i.e. near the black hole. At the same time, we know that between the black hole and the disk there are also so-called amazing S-stars [^1] Let us remind from where the fairly young S-stars (Fig. 8a in (V. D. Rusov et al., 2021)) appear in our galaxy between the black hole and the disk. Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence (S. Takekawa et al., 2019)). If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig. 8b in V. D. Rusov et al. (2021)), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig. 8b in V. D. Rusov et al. (2021)), simply flies out of the galaxy. From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., which have a little more than 30 stars. The most interesting thing is that the closest S-star to the black hole is called S-102, which has an 11-yer period of around the black hole (see Fig. 6 and Fig. 8 in V. D. Rusov et al. (2021)). From here, we understand that when the period of S-102 is very close to the black hole, then the high speed of this star is such that the dark matter is practically not captured by this star. But when the elliptical star S-102 rotates farther from the black hole, then the star gravitationally captures these ADMs.
This means that the 11-year variations in the density of dark matter via vertical density waves from the black hole disk to the solar neighborhood V. D. Rusov et al. (2021) are also captured in the solar interior! Hence, the main result is that the 11-year variations in e.g. sunspots are never associated with any dynamo action.
We could raise a few more questions, but it is enough for us that “…not enough for a scientific understanding of the solar dynamo. For there is an unfortunate revelation: with four free parameters you can provide a satisfactory fit to the horizon of New York or Beijing” – as was said by the great E. N. Parker (2009).

### v3

# A phenomenological model of magnetic flux tubes in strong fields induced by axion-origin photons

@@ Abstract
We propose a phenomenological model in which empty flux tubes in strong magnetic fields are associated with both the conversion of axions to photons of axion origin to generate coronal heating, and the simultaneous preservation of the Parker-Biermann cooling effect by converting high-energy photons from the radiative zone to axions from the convective zone through the O-loop tachocline, is the source of the rise of the magnetic flux tube from the tachocline to the photosphere in the form of sunspots.
The model is characterized by two free parameters, justified by existing theoretical and observational constraints, the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, and asymmetric dark matter (ADM) with a particle mass $m_{ADM}\sim 5~GeV$ gravitationally captured by the Sun. In this scenario, temporal modulations of the axion density are caused by anticorrelated 11-year modulations of ADM gravitationally captured by the Sun.
We also obtained theoretical estimates of the velocities and times in empty magnetic tubes in strong fields rising from the tachocline to the surface of the photosphere in the form of sunspots. This ensured complete agreement with known experimentally measured velocities and times using local helioseismological time-distance inversions along the ray trajectories of acoustic waves. This allows us to study not only the subsurface structure and dynamics of active regions, but also the depths of the empty magnetic flux tube, extending up to 200,000 km from the tachocline to the photosphere.
The proposed model links the phenomenology of axions with the magnetic structures of the Sun and provides testable consequences for astrophysical and solar observations.
@@ I Introduction
It is well known that helioseismological methods are well-suited for studying the Sun’s rotation and structure from helioseismic reversals, but they have difficulty distinguishing the effects of very strong magnetic fields from other structural changes, and therefore cannot definitively confirm the presence of a magnetic field in the solar tachocline exceeding $10^{5}~G$. In other words, it is difficult to distinguish the effects of very strong magnetic fields in the tachocline, but this does not mean they should be discarded due to their virtually undetectable nature using direct seismic methods.
This raises the question of why the dynamo action is only associated with a magnetic field of about $10^{5}~G$, but is definitely not associated with a stronger magnetic field of about $10^{7}~G$? It is known that “…even if complete knowledge of the solar dynamo existed – and it does not…”, as has been pointed out throughout the review [1] – a number of crucial questions require answers, including, at least: How do sunspots with structure, rotation velocity, and 11-year luminosity form and disappear? Why has spot formation not been captured even in local models, let alone global models? How do turbulence coefficients – turbulent pumping, turbulent diffusion – change with rotation velocity, luminosity, and internal structure? Why the inability to model accurately enough the surface layers where the density drops vertiginously is, according to [2], the common characteristic of all existing dynamo models? At that, using the so-called new era of 3D modeling and observations [1], they (solar physicists), surprisingly, believe that the study of the solar dynamo will be definitely solved within the next 10-20 years.
But the most important thing is that any dynamo action that involves, for example, the 11-year oscillations of sunspots or a sunspot butterfly diagram, must contain a small number of “free parameters” – usually three or less (see Dahlquist’s key theorem [3]) – because of the mathematical structure of the problem, where the independence of the equations, the stability of the system, and the dimensionality restrictions reduce the number of “degrees of freedom” needed to determine the solution.
We could raise a few more questions, but it is enough for us that “…not enough for a scientific understanding of the solar dynamo. For there is an unfortunate revelation: with four free parameters you can provide a satisfactory fit to the horizon of New York or Beijing” – as was said by the great Eugene N. Parker, who in 2009 in his last article [4] first began to strongly doubt the existence of the solar dynamo.
Now, in contrast to the no more than of three “free” parameters of the solar dynamo, consider our next two “free” parameters: the dark matter axions in the Sun’s core and the dark matter asymmetry (ADM) gravitationally captured by the Sun. These variables in the corresponding models cannot be predicted by the theory itself and must be determined from experimental data, as they are related to fundamental physical constants.
Let us first consider the first “free parameter” – the dark matter axions in the Sun’s core, which are necessary to solve the well-known problem of energy transfer by magnetic flux tubes from the tachocline to the Sun’s photosphere.
@@
It is known that the unsolved problem of energy transport by magnetic flux tubes at the same time represents another unsolved problem related to the sunspot darkness (see 2.2 in [5]). Of all the known concepts playing a noticeable role in understanding the connection between the energy transport and sunspot darkness, let us consider the most significant theory, in our view. It is based on the Parker-Biermann cooling effect [6, 7, 8] and originates from the early works of Biermann [7] and Alfvén [9].
@@
As you know, the Parker-Biermann cooling effect [6, 7, 8], which plays a role in our current understanding, originates from Biermann [7] and Alfvén [9]: in a highly ionized plasma, the electrical conductivity can be so large that the magnetic fields are frozen into the plasma. Biermann realized that the magnetic field in the spots themselves can be the cause of their coolness – it is colder because the magnetic field suppresses the convective heat transport. Hence, the darkness of the spot is due to a decrease in surface brightness.
@@
Parker [6, 10, 11, 12, 8] pointed out that the magnetic field can be compressed to the enormous intensity only by reducing the gas pressure within the flux tube relative to the pressure outside, so that the external pressure compresses the field. The only known mechanism for reducing the internal pressure sufficiently is a reduction of the internal temperature over several scale heights so that the gravitational field of the Sun pulls the gas down out of the tube (as described by the known barometric law $dp/dz=-\rho g$). Hence it appears that the intense magnetic field of the sunspot is a direct consequence of the observed reduced temperature [6].
@@
On the other hand, Parker [12, 13] has also pointed out that the magnetic inhibition of convective heat transport beneath the sunspot, with the associated heat accumulation below, raises the temperature in the lower part of the field. The barometric equilibrium leads to enhanced gas pressure upward along the magnetic field, causing the field to disperse rather than intensify. Consequently, [12] argued that the temperature of the gas must be influenced by something more than the inhibition of heat transport!
@@
Our idea is that the explanation of sunspots is based not only on the suppression of convective heat transfer by a strong magnetic field (of the order of $\sim 10^{7}~G$ [14]) through the enhanced cooling of the Parker-Biermann effect [10], but also on the appearance of the axions of photonic origin (Fig 2, Fig. B.1 in [14]) from the tachocline to the photosphere, which is confirmed by the “disappearance” of the heat, and consequently, the temperature in the lower part of the magnetic tube [12, 13] due to the axions of photonic origin from the photon-axion oscillations in the O-loop near the tachocline (see Fig. 2).
@@
This means that the appearance of axions of photonic origin, with which the problem of temperature increase in the lower part of a strong magnetic tube disappears, and the photons of axionic origin, which have the free path (Rosseland length; see Fig. B.3 in [14]) from the tachocline to the photosphere, are the explanation of sunspots based not only on the suppression of convective heat transport by a strong magnetic field [14], but also on the indispensable existence of the Parker-Biermann cooling effect. At the same time, we clearly understand that stronger fields (e.g. of the order of $10^{7}~G$) would seriously suppress the action of the dynamo (see [15]).
@@
On the other hand, we understand that the existence of the Parker-Biermann cooling effect is associated with the so-called thermomagnetic Ettingshausen-Nernst effect (see Apendix A in [14]). Due to the large temperature gradient in the tachocline, the thermomagnetic EN effect [16, 17, 18, 19] creates electric currents that are inversely proportional to the strong magnetic field of the tachocline. It means that, using the thermomagnetic EN effect, a simple estimate of the magnetic pressure of an ideal gas in the tachocline of e.g. the Sun,
@@
Using the strong MFT field, we are interested in the existence of dark matter axions identical to solar axions, which strongly affects the conversion of axion origin photons in magnetic O-loops near the tachocline, and are thus related to the so-called thermomagnetic Ettingshausen-Nernst effect through the Parker-Biermann cooling effect (see Apendix A in [14]). In order to answer this question, let us first consider all the unexpected and intriguing implications of the 11-year modulations of the ADM density in the solar interior and around the BH (see Sect. 3 in [14]).
It is well known that virtually all physicists, and astrophysicists in particular, know that asymmetric dark matter (ADM) exists in the Universe, and of course, in our galaxy, with the highest density of ADM located precisely at the galactic center, i.e. near the black hole. At the same time, we know that the periods, velocities, and modulations of the S-stars are the essential indicator of the modulation of the ADM halo density in the fundamental plane of the Galaxy’s center. Surprisingly, between the black hole and the disk, there are also so-called amazing S-stars  [^1] Let us remind from where the fairly young S-stars (Fig. 8a in [14]) appear in our galaxy between the black hole and the disk. Since we know that in our galaxy, in addition to the supermassive black hole (SMBH), there is the so-called intermediate black hole (IMBH; see evidence [136]). If we take into account the evolution of binary SMBH-IMBH, then when one body (SMBH) exchanges partners of binary stars orbiting the IMBH (Fig. 8b in [14]), and through the extreme gravitational-tidal field, one star is captured (SMBH) and loses energy, while the other escapes, receives all this energy and, due to the hypervelocity of B-type stars – the IMBH sling (Fig. 8b in [14]), simply flies out of the galaxy. From here it becomes clear how in other galaxies a young S-star can end up between a black hole and a disk., of which there are just over 30. Most intriguingly, the closest S-star to the black hole is called S-102, whose orbital period around the black hole is 11 years (see Fig. 6 and Fig. 8 in [14]). From this we understand that when the S-102 passes very close to the black hole, the high speed of this star is such that dark matter is practically not captured by this star. But when the star S-102 orbits farther from the black hole, the star gravitationally captures these ADMs. If the modulations of the ADM halo at the GC lead to modulations of the ADM density on the surface of the Sun (through vertical density waves from the disk to the solar neighborhood), then there is an ”experimental” anticorrelation identity between such indicators as the ADM density modulation in the solar interior and the sunspot cycles. Or equivalently, between the modulation of solar axions (or photons of axion origin) and the sunspot cycles!
@@
A hypothetical pseudoscalar particle called axion is predicted by the theory related to solving the CP-invariance violation problem in QCD. The most important parameter determining the axion properties is the energy scale $f_{a}$ of the so-called U(1) Peccei-Quinn symmetry violation. It determines both the axion mass and the strength of its coupling to fermions and gauge bosons including photons. However, in spite of the numerous direct experiments, axions have not been discovered so far. Meanwhile, these experiments together with the astrophysical and cosmological limitations leave a rather narrow band for the permissible parameters of invisible axion (e.g. $10^{-6}eV\leqslant m_{a}\leqslant 10^{-2}eV$ (see [21, 22]; Fig. 13 in [23], Fig. 24 in [24])). The PQ mechanism, solving the strong CP problem in a very elegant way [25, 26, 27, 28], is especially attractive here, since the axion is also a candidate for dark matter [29, 30, 31].
@@
At present, we have shown that axions born in the core of the Sun, and photons of axion origin (in the million-degree solar corona) yield a tight constraint on the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$ and mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ (see Fig. 1a; also Fig.1 in [14]). Most importantly, hadronic axions produced in the solar core are controlled by anticorrelated 11-year variations in the density of asymmetric dark matter (ADM) gravitationally trapped in the solar interior (via vertical density waves from the black hole disk to the solar neighborhood), with the hard part of the solar photon spectrum being of axion origin, which is defined as the product of the axions fraction in the solar core and the fraction of the sunspot area. The latter is the result of solar luminosity variations (see Eqs. (27)-(28) in [14]). This means that solar luminosity variations and other solar cycles are the result of 11-year variations in ADM density, which are anticorrelated with the density of hadronic axions in the solar core.
@@
Interestingly, in contrast to our Fig. 1a, in Fig. 1b, we can see a well-known strong constraint on the axion-photon coupling $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, but in a wide mass range (see Fig. 1b and Eq. (2) in [32]) from an analysis of a sample of 39 Galactic Globular Clusters. However, in our case (Fig. 1a) the wide range of masses [32] is, oddly enough, replaced by a single mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$, which fits not only the axions of galactic globular clusters [32], but also all axion stars in our galaxy, including neutron stars [33], supernovae [34, 35], and the Sun [14].
Figure 1: (a) Summary of astrophysical, cosmological and laboratory constraints on axions. Comprehensive axion parameter space, highlighting two main front lines of direct detection experiments: helioscopes (CAST [36]) and haloscopes (ADMX [37], and RBF [38] and microwave resonators [39, 40, 41, 42, 43]). The astrophysical bounds from horizontal branch and massive stars are labeled “HB” [44] and “Cepheids” [45], respectively, and there are also astrophysical hints (WD cooling hints, and HB hint). The QCD motivated models (KSVZ [46, 47]) for axions lay in the yellow diagonal band. A plot of $g_{a\gamma}$ versus $m_{a}$ with the most stringent results (solid lines) and sensitivity perspectives (dashed lines) of observations and experiments directly comparable to the different phases of IAXO are shown, BabyIAXO, IAXO, and an upgraded version of IAXO, IAXO+. The yellow band denotes the region of the parameter space favoured by QCD axion models. The red star marks the values of the axion mass $m_{a}\sim 3.2\cdot 10^{-2}~eV$ and the axion-photon coupling constant $g_{a\gamma}\sim 4.4\cdot 10^{-11}~GeV^{-1}$, which were first obtained experimentally (see [14]). (b) $R$ parameter constraints, which compares the numbers of stars in the the horizontal branch (HB) and in the upper portion of the red giant branch (RGB), to helium mass fraction $Y$ and axion coupling $g_{a\gamma}$ (adopted from [32]). The resulting bound on the axion ($g_{10}=g_{a\gamma}/(10^{-10}~GeV^{-1})$ is somewhere between rather conservative $0.5<g_{10}<0.8$ and most aggressive $0.35<g_{10}<0.5$ [48]. The red line marks the value of the axion–photon coupling constant $g_{a\gamma}\sim 4.5\cdot 10^{-11}~GeV^{-1}$ adopted from Eq. (2) in [32]. The blue shaded area represents the bounds from Cepheids observation. The yellow star corresponds to $Y=0.254$ and the bounds from HB lifetime (yellow dashed line).
@@
II Physics of the theory of practically empty flux tubes in strong fields of the order $\sim 10^{7}~G$